the roulette has 43 pockets (20 red, 20 black, and three green pockets). assume you lose the whole bet if the ball falls into a pocket with a different color. 1) how much money do you lose in the long run per game if you always bet 10 euro on the green color playing version y? (win: +100 euro | lose: -10 euro) (1p.)

1
Expert's answer
2021-11-16T16:21:47-0500

Chance of green color falling out "\frac{3}{43}", chance of non-falling out "\frac{40}{43}". Calculate the mathematical expectation:

"E=\frac{3}{43}\cdot100-\frac{40}{43}\cdot 10=-\frac{100}{43}"

Answer: "\frac{100}{43}" euro will lose in the long run.